# Anomaly Detection in Complex Dynamical Systems: A Systematic Framework Using Embedding Theory and Physics-Inspired Consistency

### v2

@@ Abstract
Anomaly detection in complex dynamical systems is essential for ensuring reliability, safety, and efficiency in industrial and cyber-physical infrastructures. Predictive maintenance helps prevent costly failures, while cybersecurity monitoring has become critical as digitized systems face growing threats. Many of these systems exhibit oscillatory behaviors and bounded motion, requiring anomaly detection methods that capture structured temporal dependencies while adhering to physical consistency principles. In this work, we propose a system-theoretic approach to anomaly detection, grounded in classical embedding theory and physics-inspired consistency principles. We build upon the Fractal Whitney Embedding Prevalence Theorem that extends traditional embedding techniques to complex system dynamics. Additionally, we introduce state-derivative pairs as an embedding strategy to capture system evolution. To enforce temporal coherence, we develop a Temporal Differential Consistency Autoencoder (TDC-AE), incorporating a TDC-Loss that aligns the approximated derivatives of latent variables with their dynamic representations. We evaluate our method on the C-MAPSS dataset, a benchmark for turbofan aeroengine degradation. TDC-AE outperforms LSTMs and Transformers while achieving a nearly 200x reduction in MAC operations, making it particularly suited for lightweight edge computing. Our findings support the hypothesis that anomalies disrupt stable system dynamics, providing a robust signal for anomaly detection.
Index Terms:
Complex Dynamical Systems, Anomaly Detection, System Theory, Embedology, Physics-Informed Machine Learning, Predictive Maintenance, Edge Computing
@@ II-A Time-Series Modeling
Models like LSTMs and Transformers with attention mechanisms have been successfully applied to time-series anomaly detection in fields such as mechanical engineered system, aerospace, and industrial monitoring [6, 7, 8]. While these methods effectively capture long-range dependencies and irregular temporal patterns, they are computationally expensive. For example, evaluated on common benchmark datasets, sequence lengths up to 500 time steps were required to achieve competitive performance [9]. The high memory requirements often exceed the constraints of typical MCU-level devices, making real-time deployment challenging. For example, a typical LSTM with 500 time steps and 50 hidden units already surpasses the available memory of common low-power devices, limiting its practical applicability. Additionally, the sequential nature of RNNs restricts parallelization, further increasing computational cost [10]. These challenges highlight the need for alternative approaches that balance computational efficiency with robust anomaly detection, aligning with broader goals of sustainability and practical deployability.
@@ II-C Benchmark Use Case: Turbofan Aeroengine Degradation
Previous work on anomaly detection in this setting has explored various deep learning models. One study employed an LSTM-based approach [19], leveraging recurrent structures to model time dependencies. Another approach used a standard autoencoder (AE) without explicit temporal modeling [18]. More recent research has investigated Transformer-based models, which excel at capturing long-range dependencies but introduce high computational costs and require extensive training data due to their large parameter space [20, 7]. Moreover, both Transformer studies formulated the problem as a multiclass classification task, with one using a slightly different dataset and the other applying the method directly to C-MAPSS. However, defining well-separated fault categories in a complex dynamical system is challenging in practice, as non-trivial interactions between multiple physical components create highly unpredictable behaviors. As system dynamics grow more intricate, these interactions become even less predictable, further complicating precise fault categorization.
II-D Proposed Approach
Instead of relying on complex models to differentiate normal and abnormal behavior, we adopt a system-theoretic approach inspired by classical embedding theory to analyze the system’s latent dimensionality. By incorporating physics-inspired consistency principles, we approximate the underlying causal mechanisms without explicitly enforcing physical laws. We hypothesize that in a stable regime, complex systems exhibit predictable behavior, enabling well-approximated low-dimensional embeddings. In contrast, anomalies introduce additional complexity, disrupting these stable relationships. We aim to leverage this property to detect abnormal states when the learned embedding no longer adequately captures the system dynamics.
@@ III Theoretical Foundation & Methods
We study the mathematical foundations of embedding bounded dynamical systems into an autoencoder’s latent space, aiming to ensure reliable mapping of system dynamics for anomaly detection and health monitoring.
Consider for example an oscillatory dynamical system, specifically a damped-driven pendulum. Initially, the system remains within a bounded oscillatory regime; however, due to the influence of drift terms, it gradually deviates, eventually departing from the confined state space. This perturbation effect is illustrated in Figure 1, where the phase space trajectory initially remains within a bounded region but progressively transitions into an unbounded state. This formulation highlights that the state variables of an oscillatory system, namely the angular displacement $\theta$ and its derivative $\dot{\theta}$, define a phase space where the system’s behavior can be systematically analyzed. Within this phase space, bounded regions correspond to normal oscillations, where the system follows a predictable, periodic trajectory. However, once the system leaves this bounded region, it signifies the onset of unbounded oscillations or deviations from regular behavior. For formal definitions and details on the system’s implementation, refer to Appendix A.
We explore whether this concept extends to complex dynamical systems and the constraints and conditions for a successful transition. The challenge lies in preserving phase space properties in high-dimensional, nonlinear systems while maintaining topological consistency and dynamical coherence. To address this, we examine the mathematical and computational requirements for embedding system dynamics into a reduced representation, guiding the design of effective autoencoder-based anomaly detection.
III-A Embedding Complex System Dynamics into a Latent Space
@@
We analyze the latent space encoding of a dynamical system from a systems-theoretic perspective. Ideally, such systems in a stable state evolve on smooth manifolds, but real-world dynamics often unfold on complex, non-smooth sets due to measurement noise and the inherent need for machines to exhibit more intricate dynamics to perform their functions effectively. A key aspect of this generalization is the notion of dimensionality for compact subsets, which leads to the definition of box-counting dimension.
Definition 1 (Box-Counting Dimension [21]).
The box-counting dimension of $A$ is defined as
provided the limit exists.
In this context, $\epsilon$ provides a measure of the complexity exhibited by motion beyond smooth manifold assumptions.
Equipped with this notion of dimensionality, we now state the embedding theorem that generalizes classical results to fractal structures.
Theorem 1 (Fractal Whitney Embedding Prevalence Theorem [22]).
  1. 1.
    $F$ is one-to-one on $A$.
  2. 2.
We discuss the necessary assumptions to apply this formalism to our system representation.
Compactness of the System State: We assume that the key drivers of the physical process form a bounded and closed subset of $\mathbb{R}^{k}$, ensuring compactness. This assumption reflects the practical reality that real-world physical systems with inherently bounded dynamics, when operating in a stable state, remain confined within a finite region of their state space.
Informative Measurement Function: We assume that the measurement function retains sufficient information about the system dynamics. Specifically, it must be injective enough over $A$ to avoid collapsing distinct states, ensuring that the measurement space captures the necessary complexity of the system.
Dimensionality of the System: We assume that the measurement-level set $A\subset\mathbb{R}^{m}$ has a well-defined box-counting dimension $d$, which reflects the effective complexity of the system’s dynamics. This dimensionality determines the necessary embedding space.
Under these conditions, the theorem is applicable. This allows us to estimate the required embedding dimension $n$ to achieve a faithful description of the system’s dynamics.
@@
After analyzing the necessary embedding dimension, we now turn to the type of embedding that best captures the dynamics of the system. We draw inspiration from simple dynamical systems and propose using state-derivative pairs as an embedding strategy. This choice is motivated by two key principles. Physical laws typically encode causal relationships between a system’s state and its derivatives, making this a natural way to describe a dynamical process. When chosen properly, state-derivative pairs help reduce redundancy by maintaining linear independence between embedding dimensions.
@@
To effectively utilize this type of embedding, we need more than just a one-to-one mapping—we also require consistent differential structure. In other words, small changes in the latent space should produce meaningful, non-degenerate changes in the measurement space. A full-rank Jacobian is a necessary and sufficient condition to guarantee this kind of well-behaved, locally reversible mapping [23].
We hypothesize that leveraging this type of embedding allows for two key insights:
  • Deviations beyond a certain region associated with normal variability indicate anomalous behavior, as suggested by phase space considerations.
  • As the dynamics become more complex, this leads to an increase in the box-counting dimension. By Th. 1, this increase suggests that the chosen embedding dimension is insufficient to fully describe the system. Consequently, we hypothesize that this insufficiency results in a collapse of the dynamical description, leading to the loss of the causal approximation between the state and its derivative.
III-B Application to Autoencoder-based Representations
@@
With the theoretical framework established, we apply it to a practical case in two steps. First, we select a stable, bounded dynamical system with accessible measurements. For this, we use the NASA C-MAPSS dataset, which provides time-series sensor data from aircraft engines. Second, we develop an autoencoder that learns embeddings aligned with our framework, ensuring a one-to-one mapping and immersion within the normal operating regime of the system.
III-B1 C-MAPSS Dataset
@@
We use the FD001 subset of the C-MAPSS dataset, which contains run-to-failure time series from 100 independent engines operating under a single operating condition but exhibiting varying fault progressions. This substantial number of individual systems allows us to assess the generalization capability of our method across multiple degradation trajectories. For anomaly detection, we follow a 60/40 split, where the first 60% of cycles are labeled as normal and the remaining 40% as anomalous, as suggested in similar studies [18, 19]. A random 80/20 train-test split based on separate engines ensures robust evaluation.
@@
III-B2 Applying the Theoretical Foundation
@@
The NASA C-MAPSS dataset provides a measurement space of dimension $k=24$, representing sensor readings from the turbine system. To determine a suitable embedding dimension in accordance with Theorem 1, we establish the following assumptions about the system’s essential dynamics.
As argued in the theory section, the turbine operates in a stable regime, where its state evolution remains within a well-defined, physically constrained region. This ensures that the system’s key drivers form a bounded and closed subset of $\mathbb{R}^{k}$, thereby satisfying the compactness assumption.
We hypothesize that the turbine’s core behavior is governed by three fundamental state-derivative pairs, corresponding to rotation, temperature, and pressure along with their respective rates of change. These represent the system’s intrinsic dynamics and serve as a basis for estimating its fractal dimension. In Appendix B, we computed the box-counting dimension numerically for the attractor of a damped driven pendulum, yielding $d\approx 1.3$. For a perfect mathematical pendulum, this value reduces to $d=1$, reflecting its simple periodic structure (a circle). Given the structural similarities between the fundamental dynamics of the pendulum and the turbine’s primary state variables, we assume a similar level of dynamic complexity here. However, this framework provides room to incorporate domain knowledge, such as machine tolerances and operational constraints, to refine this assumption further. This leads to an estimated box-counting dimension $d<4$ for the combined system. Applying Theorem 1, this result suggests the selection of an embedding dimension of $n=8$.
The measurement function in the C-MAPSS dataset is assumed to retain sufficient information about the system’s essential dynamics. Given the structured nature of the sensor readings, it provides a reliable representation of the turbine’s state evolution without collapsing distinct states, thereby satisfying the injectivity condition.
Based on our approach, we select four state-derivative pairs for the embedding. To ensure that this representation is at an immersion in the subspace corresponding to the system’s normal or stable regime, we apply a numerical method to verify whether the Jacobian has full rank in the normal state of the test data. While this does not constitute a strict mathematical proof, it serves as a reliable guideline for assessing the validity of the embedding. SVD is particularly powerful for analyzing the rank of a matrix. The rank of $X$ can be determined by counting the number of non-zero singular values in $\Sigma$. If all singular values are non-zero, the matrix is full rank [28].
@@
Theorem 1 assumes a smooth map and that it applies to ”almost every” map. A $\tanh$-activated neural network encoder reasonably satisfies the smoothness assumption. The ”almost every” condition applies to all smooth maps, whereas neural networks form a restricted subset. However, by universal approximation, a sufficiently large network can approximate such an embedding, making the theorem a strong and relevant practical guideline.
III-B3 Temporal Differential Consistency informed Autoencoder
@@
As a foundational step, we require a method to approximate the first-time derivative. A widely used approach is the central difference method, which estimates the derivative of a function by computing the slope between two points symmetrically positioned around the point of interest. The central difference formula for the first derivative of a function $f(t)$ is:
$\dot{f}(t)\approx\frac{f(t+\Delta t)-f(t-\Delta t)}{2\Delta t}$
where $\Delta t$ represents a small step size. This method is often preferred over forward or backward differences due to its higher accuracy ($O(\Delta t^{2})$), meaning that the approximation error decreases quadratically as $\Delta t$ decreases [29].
@@
To accurately capture temporal dynamics, we incorporate the central difference method into the training framework of the latent space. The latent representations at the previous ($t-1$) and next ($t+1$) time steps are used to approximate the first-time derivative of the static latent variables ($z$) using the central difference formula. This derivative is then used as a target for the dynamic latent variables ($\dot{z}$), ensuring consistency with the central difference approximation, scaled by the time interval $\Delta t$. A schematic representation of this approach is shown in Fig. 7 in Appendix C.
@@
III-B4 Consistency Metrics
We employ an embedding based on state-derivative pairs, where the derivative is approximated using the central difference method. While numerical differentiation is generally ill-posed, particularly in the presence of noise [30], our approach does not seek a fully precise description of the system’s dynamics. Instead, we aim for a causal approximation that reliably indicates when the system deviates from normal variability. By choosing a sufficiently large embedding dimension, we ensure that the system operates on a simple attractor geometry. This allows us to expect a nearly constant trendline for both the state and its derivative. To evaluate this, we introduce two metrics

### v3

@@ Abstract
Anomaly detection in complex dynamical systems is essential for ensuring reliability, safety, and efficiency in industrial and cyber-physical infrastructures. Predictive maintenance helps prevent costly failures, while cybersecurity monitoring has become critical as digitized systems face growing threats. Many of these systems exhibit oscillatory behaviors and bounded motion, requiring anomaly detection methods that capture structured temporal dependencies while adhering to physical consistency principles. In this work, we propose a system-theoretic approach to anomaly detection, grounded in classical embedding theory and physics-inspired consistency principles. We build upon the Fractal Whitney Embedding Prevalence Theorem that extends traditional embedding techniques to complex system dynamics. Additionally, we introduce state-derivative pairs as an embedding strategy to capture system evolution. To enforce temporal coherence, we develop a Temporal Differential Consistency Autoencoder (TDC-AE), incorporating a TDC-Loss that aligns the approximated derivatives of latent variables with their dynamic representations. We evaluate our method on two subsets (FD001, FD003) of the C-MAPSS dataset, a benchmark for turbofan engine degradation. TDC-AE machtes LSTMs and outperforms Transformers while achieving a nearly 100x reduction in MAC operations, making it particularly suited for lightweight edge computing. Our findings support the hypothesis that anomalies disrupt stable system dynamics, providing a robust signal for anomaly detection.
@@ II-A Time-Series Modeling
Models like LSTMs and Transformers have been successfully applied to time-series anomaly detection in fields such as mechanical engineered system, aerospace, and industrial monitoring [6, 7, 8]. While these methods effectively capture long-range dependencies and irregular temporal patterns, they are computationally expensive. For example, evaluated on common benchmark datasets, sequence lengths up to 500 time steps were required to achieve competitive performance [9]. The high memory requirements often exceed the constraints of typical MCU-level devices, making real-time deployment challenging. For example, a typical LSTM with 500 time steps and 50 hidden units already surpasses the available memory of common low-power devices, limiting its practical applicability. Additionally, the sequential nature of RNNs restricts parallelization, further increasing computational cost [10]. These challenges highlight the need for alternative approaches that balance computational efficiency with robust anomaly detection, aligning with broader goals of sustainability and practical deployability.
@@ II-C Benchmark Use Case: Turbofan Aeroengine Degradation
Previous work on anomaly detection in this setting has explored various deep learning models. One study employed an LSTM-based approach [19], leveraging recurrent structures to model time dependencies. Another approach used a Vanilla Autoencoder (AE) without explicit temporal modeling [18]. More recent research has investigated Transformer-based models, which excel at capturing long-range dependencies but introduce high computational costs and require extensive training data due to their large parameter space [20, 7]. Moreover, both Transformer studies formulated the problem as a multiclass classification task, with one using a slightly different dataset and the other applying the method directly to C-MAPSS. However, defining well-separated fault categories in a complex dynamical system is challenging in practice, as non-trivial interactions between multiple physical components create highly unpredictable behaviors. As system dynamics grow more intricate, these interactions become even less predictable, further complicating precise fault categorization.
@@ III Theoretical Foundation & Methods
III-A Proposed Approach & Motivation
We propose a data-driven anomaly detection framework grounded in physics consistency and embedding theory, without explicitly enforcing known physical laws. Our central premise is that effective anomaly detection in dynamical systems requires understanding their intrinsic dynamics.
We model system behavior using state-derivative pairs, a natural representation for dynamical systems. From observed measurements, we construct a mapping into a state-space-like manifold that approximates the system’s latent dynamics. Under normal operating conditions, we hypothesize that the system evolves within a well-defined, compact set of dynamic states, and that anomalies manifest as deviations from this set.
Consider, for example, a damped-driven pendulum. Its phase space, defined by angular displacement $\theta$ and its derivative $\dot{\theta}$, initially confines the system to a bounded, oscillatory regime. However, under drift, the trajectory may exit this region, signaling a departure from regular dynamics. This behavior exemplifies how anomalies manifest as transitions from compact, predictable trajectories to unbounded or inconsistent dynamics (see Fig.1, implementation details in Appendix A).
To justify this assumption, the learned mapping must be locally invertible and injective in the normal regime, ensuring that nearby observations correspond to unique and consistent states in the latent space.
Embedding theory provides guidance for achieving injectivity by relating the geometric complexity of the measurement manifold to a lower bound on the embedding dimension required for faithful reconstruction of system dynamics.
We assume that anomalies introduce additional dynamic complexity, which increases the effective geometric complexity of the measurement space. Our goal is to construct an embedding whose geometric complexity is matched to the dynamic complexity of normal operation, sufficient to represent typical behavior, but not overparameterized to absorb arbitrary deviations.
When an anomaly occurs, the resulting increase in complexity exceeds the representational capacity of the embedding, leading to an inconsistent or distorted description of the latent dynamics. We exploit this mismatch as a signal for anomaly detection.
To validate our assumptions, we use numerical methods to verify that the Jacobian of the learned mapping is full-rank (for invertibility) and to empirically assess local injectivity.
III-B Embedding Complex System Dynamics into a Latent Space
@@
Embedding measurements into a lower-dimensional space, as described above, is a well-studied problem in classical embedology. Embedding theorems, such as those by Whitney, provide conditions under which one can construct injective maps from measurement data to a lower-dimensional space. Specifically, these theorems relate the dimensionality of the manifold covered by the measurements to the ambient space in which the measurements reside.
Ideally, systems in stable operation evolve on smooth manifolds. However, real-world dynamics often lie on more complex, non-smooth sets due to measurement noise and the inherently intricate behaviors required by physical systems.
To generalize the idea of dimension for such cases, we use the box-counting dimension, which extends naturally to non-smooth subsets. For a heuristic and accessible explanation of this concept and a formal definition, see Appendix A.)
The Fractal Whitney Embedding Prevalence Theorem [21] (see Appendix A) states that if the measurement data lie in a compact subset of the measurement space, and if the box-counting dimension of this set is $d$, then an embedding into $\mathbb{R}^{n}$ with $n>2d$ is generically injective. This provides a practical criterion: estimate the box-counting dimension of the measurements and choose an embedding dimension that exceeds twice this value to obtain an injective mapping with high probability.
@@
Having established the role of embedding theory, we now turn to physical consistency part. We draw inspiration from simple dynamical systems and propose using state-derivative pairs as the embedding. This choice is motivated by two principles. First, physical laws typically encode causal relationships between a system’s state and its derivative, making this representation natural for dynamical processes. Second, state-derivative pairs help reduce redundancy by promoting linear independence between embedding dimensions.
@@
To be effective, such embeddings must not only be injective, but also preserve a consistent differential structure. That is, small changes in the latent space should induce meaningful, non-degenerate changes in the measurement space. A full-rank Jacobian is a necessary and sufficient condition to ensure this well-behaved, locally reversible mapping [22].
We use intrinsic dimensionality (ID) as a proxy for the boxcounting dimension. We applied a per-engine estimation procedure using the Two-Nearest-Neighbor (Two-NN) method [23]. This estimator is particularly well suited for our setting, where each engine provides a small number of samples that are assumed to lie on a low-dimensional manifold.
To empirically assess whether our mapping preserves injectivity, we compute the ratio of pairwise distances in the latent space $Z$ to those in the input space $X$, i.e., $\|\mathcal{E}(x_{1})-\mathcal{E}(x_{2})\|/\|x_{1}-x_{2}\|$, for input pairs satisfying $\|x_{1}-x_{2}\|>\delta$ (with $\delta=10^{-6}$). A minimum ratio below $10^{-5}$ indicates potential collapse and loss of injectivity. This threshold is chosen based on standard floating-point precision limits and empirical practices in representation learning [24]. The method serves as a practical diagnostic for identifying when distinct inputs are mapped to nearly identical representations.
III-C Application to Autoencoder-based Representations
@@
With the theoretical framework established, we select a stable, bounded dynamical system with accessible measurements. For this, we use the NASA C-MAPSS dataset, which provides time-series sensor data from aircraft engines.
III-C1 C-MAPSS Dataset
@@
We use the FD001 and FD003 subsets of the C-MAPSS dataset, which differ in fault complexity: FD001 includes one fault mode, while FD003 includes two. For anomaly detection, we follow a 60/40 split, where the first 60% of cycles are labeled as normal and the remaining 40% as anomalous, as suggested in similar studies [18, 19]. A random 80/20 train-test split based on separate engines ensures robust evaluation.
@@
III-C2 Applying the Theoretical Foundation
@@
The NASA C-MAPSS dataset provides a measurement space of dimension $k=24$, representing sensor readings from the turbine system. To determine a suitable embedding dimension in accordance with Th. 1, we establish the following assumptions about the system’s essential dynamics.
We assume that the measurement function in the C-MAPSS dataset retains sufficient information about the underlying system dynamics. Given the structured and high-resolution nature of the sensor data, the measurement space is taken to provide a reliable, non-degenerate representation of the turbine’s evolving state—a standard assumption in physical modeling from observation data.
Intrinsic Dimensionality and Embedding.
Using the Two-NN estimator as a proxy for the box-counting dimension, we obtain the following intrinsic dimension estimates for each dataset subset: F001: $4.93\pm 0.75$, and F003: $4.83\pm 0.65$. Interpreting these as fractal dimensions of the measurement space, and applying Theorem 1, we require an embedding space of dimension greater than twice the fractal dimension. This yields lower bounds of $9.86$ and $9.66$, respectively. Rounding to the next even integer to accommodate both state and state-derivative pairs, we select embedding dimensions $n=10$ for both F001 and F003.
@@
Th. 1 assumes a smooth map and that it applies to ”almost every” map. A $\tanh$-activated neural network encoder reasonably satisfies the smoothness assumption. The ”almost every” condition applies to all smooth maps, whereas neural networks form a restricted subset. However, by universal approximation, a sufficiently large network can approximate such an embedding, making the theorem a strong and relevant practical guideline.
Rank Verification via SVD.
To ensure that the reconstructed state matrix $X$ captures the full dynamics, we verify its rank using singular value decomposition (SVD). The rank is determined by the number of non-zero singular values in $\Sigma$; if all are non-zero, the matrix is full-rank [29].
III-C3 Temporal Differential Consistency informed Autoencoder
@@
As a foundational step, we require a method to approximate the first-time derivative. A widely used approach is the central difference method, which estimates the derivative of a function by computing the slope between two points symmetrically positioned around the point of interest. This method is often preferred over forward or backward differences due to its higher accuracy ($O(\Delta t^{2})$), meaning that the approximation error decreases quadratically as $\Delta t$ decreases [30].
@@
To accurately capture temporal dynamics, we incorporate the central difference method into the training framework of the latent space. The latent representations at the previous (t1t-1) and next ($t+1$) time steps are used to approximate the first-time derivative of the static latent variables ($z$) using the central difference formula. This derivative is then used as a target for the dynamic latent variables ($\dot{z}$), ensuring consistency with the central difference approximation, scaled by the time interval $\Delta t$. A schematic representation of this approach is shown in Fig. 7 in Appendix C.
@@
III-C4 Consistency Metrics
We employ an embedding based on state-derivative pairs, where the derivative is approximated using the central difference method. While numerical differentiation is generally ill-posed, particularly in the presence of noise [31], our approach does not seek a fully precise description of the system’s dynamics. Instead, we aim for a causal approximation that reliably indicates when the system deviates from normal variability. By choosing a sufficiently large embedding dimension, we ensure that the system operates on a simple attractor geometry. This allows us to expect a nearly constant trendline for both the state and its derivative. To evaluate this, we introduce two metrics, which are detailed in the following paragraphs.